I am having some trouble with this question so if could you please help me with this probelm it would be greatly appricated. Thank you very much. Here is the question...

Complementary angles are two angles whose measures add to 90degrees. The ratio of the measures of two complementary angles is 4:11. What are the measures of the angles?

Answer by uma(370)


Given that the ratio of the measures of two complementary angles is 4:11
Let us take the angles to be 4x and 11x as the ratio is 4:11
They are complementary implies, 4x + 11x = 90
==> 15x = 90
==> x = 90/15
==> x = 6
Thus the angles are 4*6 = 24 and 11*6 = 66
Answer : 24 degrees and 66 degrees
Good Luck!!!